# Complete interpolating sequences for Fock type spaces

### 16 Mar 2026 v1

@@ 5. Separated sequences
One of the main tools we rely on is the Bernstein-type inequality. We begin with the following lemma,
@@
$\left|\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})}\right|^{p}$ $\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-\ell(y_{n})}e^{\ell(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})^{p}}{|z-\sigma_{n}|^{p}}.$ (5.1)
@@
$\frac{\left|g_{\sigma_{n}}(z)\right|^{p}}{\|L_{z}\|^{p}}\asymp e^{-\ell(y_{n})+\ell(t)-2t-pnt+\psi^{\prime}(y_{n})y_{n}},\quad z\in D(\sigma_{n},\varepsilon|\sigma_{n}|).$ (5.2)
@@
$|a_{m,n}|^{p}\asymp\frac{(\psi^{\prime\prime}(y_{n}))^{1/2}}{(\psi^{\prime\prime}(y_{s_{m}}))^{1/2}}e^{\vartheta(m,n)}\frac{|\gamma_{m}-\sigma_{s_{m}}|^{p}}{|\gamma_{m}-\sigma_{n}|^{p}}$ (5.3)
@@
This concludes the proof. ∎
@@
$\lesssim|z-w|\sum_{n}|a_{n}|\frac{(\psi^{\prime\prime}(y_{n}))^{\frac{1}{2p}}}{|z-\sigma_{n}||w-\sigma_{n}|}e^{\frac{1}{p}(\psi(y_{n})-\ell(y_{n}))}..$ (5.4)
$\bullet$ The case where $1<p<\infty$. Let $q$ be the Hölder conjugate of $p$. Using Hölder’s inequality we obtain
@@
If $\Gamma$ is an interpolating sequence for ${\mathcal{F}}^{p}_{\varphi}$, $0<p<\infty$, then for every $k$ there exists $f_{k}\in\mathcal{F}^{p}_{\varphi}$ such that
@@
Hence $d_{\log}(\gamma_{k},\gamma_{j})\gtrsim 1$ and hence $\Gamma$ is logarithmically separated. The proof for the case $p=\infty$ is similar and is therefore omitted. ∎
@@ 6.1. Sufficient conditions
$f_{v}(z):=\sum_{n}v_{n}\|L_{\gamma_{n}}\|\frac{G_{\Gamma}(z)}{G^{\prime}_{\Gamma}(\gamma_{n})(z-\gamma_{n})},\quad z\in\mathbb{C}.$ (6.1)
@@
$A_{n,m}:=\frac{\|L_{\gamma_{n}}\|}{\|L_{\sigma_{m}}\|}\frac{G_{\Gamma}(\sigma_{m})}{G^{\prime}_{\Gamma}(\gamma_{n})(\sigma_{m}-\gamma_{n})}$ (6.2)
@@
$\left|A_{n,m}\right|\lesssim\exp\left(-\left[\frac{1}{2}-\Delta_{N}-\varepsilon\right]|y_{m+j}-y_{n}|\right).$ (6.3)
@@ 6.2. Necessary conditions
$|f_{\gamma}(z)|^{p}\lesssim\|f_{\gamma}\|^{p}_{\varphi,p}\|L_{z}\|^{p}\lesssim\frac{\left(\psi^{\prime\prime}(\log|z|)\right)^{\frac{1}{2}}}{\|L_{\gamma}\|^{p}(1+|z|^{2})}e^{p\varphi(z)},\quad z\in\mathbb{C}.$ (6.4)
@@
$\left|G_{\Gamma}(\sigma_{m_{k}})\right|\asymp\frac{|\sigma_{m_{k}}-\gamma_{n_{k}}|}{|\gamma_{n_{k}}|}\prod_{j=0}^{n_{k}-1}\frac{\sigma_{m_{k}}}{|\gamma_{j}|},\qquad\left|G^{\prime}_{\Gamma}(\gamma_{n_{k}})\right|\asymp\frac{1}{|\gamma_{n_{k}}|}\prod_{j=0}^{n_{k}-1}\frac{|\gamma_{n_{k}}|}{|\gamma_{j}|}.$ (6.5)
@@
$e^{pn_{k}(y_{m_{k}}-x_{n_{k}})}=\left(\frac{\sigma_{m_{k}}}{|\gamma_{n_{k}}|}\right)^{pn_{k}}\asymp\left|f_{\gamma_{n_{k}}}(\sigma_{m_{k}})\right|^{p}\lesssim\frac{\left(\psi^{\prime\prime}(y_{m_{k}})\right)^{\frac{1}{2}}}{\|L_{\gamma_{n_{k}}}\|^{p}}e^{\psi(y_{m_{k}})-2y_{m_{k}}}.$ (6.6)
@@
$A_{n,m}:=\frac{\|L_{\gamma_{n}}\|}{\|L_{\sigma_{m}}\|}\frac{G_{\Gamma}(\sigma_{m})}{G^{\prime}_{\Gamma}(\gamma_{n})(\sigma_{m}-\gamma_{n})}$ (6.7)
@@
$p\sum_{k=1}^{m}y_{k}=-\psi^{\prime}(y_{m})y_{m}+\psi(y_{m})-\frac{p}{2}y_{m}-\left(\frac{p^{2}}{12}+o(1)\right)\frac{1}{\psi^{\prime\prime}(y_{m})}+O(1).$ (6.8)
@@
$|A_{n,m}|^{p}\gtrsim\exp\left[-\frac{p}{2}|y_{n}-y_{m+j}|+C_{\psi}\left(\frac{1}{\psi^{\prime\prime}(y_{n})}-\frac{1}{\psi^{\prime\prime}(y_{m+j})}\right)\pm p\sum_{k=m+j+1}^{n}\delta_{k}\right].$ (6.9)
@@
$F_{v}(z)=\sum_{n\geq-1}v_{n}e^{\varphi(\gamma_{n})}\frac{G_{\Gamma^{*}}(z)}{G_{\Gamma^{*}}^{\prime}(\gamma_{n})(z-\gamma_{n})}$ (7.1)
@@
To prove Theorem 1.3, we adopt an approach inspired by [4, 23]. To establish the necessary conditions for a sequence to be sampling or interpolating, we rely on the following lemma. The proof in our setting is identical to that of [4, Lemma 5.1] (see also [20, Lemma 40]), using the point evaluation norm estimates and the canonical product bounds derived above.
@@
$\frac{1}{y_{(kN+2)M}-y_{(kN+1)M}}\left|\sum_{n=(kN+1)M+1}^{(kN+N+1)M}\ \delta_{n}\right|\ \leq\ C.$ (8.1)
@@
$|\delta_{n}|<\eta(pM(m+1)+2-\frac{p}{2})-\eta(pmM+2-\frac{p}{2})$ (8.2)
@@
$|\delta_{n}|<\eta(pM(j+1)+2-\frac{p}{2})-\eta(pjM+2-\frac{p}{2})$ (8.5)

### 23 Mar 2026 v2

@@ 5. Separated sequences
A sequence $\Gamma$ is said to be logarithmically separated if there exists a constant $d_{\Gamma}>0$ such that
$\inf\{d_{\log}(\gamma,\gamma^{*}):\gamma,\gamma^{*}\in\Gamma,\ \gamma\neq\gamma^{*}\}\geq d_{\Gamma},$
where
$d_{\log}(z,w)=\frac{|z-w|}{1+\min(|z|,|w|)},\qquad z,w\in\mathbb{C}.$
One of the main tools we rely on is the following Bernstein-type result (see Lemma 5.2). We begin with the following lemma, which shows that the inequality below allows us to prove that the sequence is a finite union of logarithmically separated sequences.
@@
(5.1) $\left|\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})}\right|^{p}$ $\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-\ell(y_{n})}e^{\ell(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})^{p}}{|z-\sigma_{n}|^{p}}.$
@@
(5.2) $\frac{\left|g_{\sigma_{n}}(z)\right|^{p}}{\|L_{z}\|^{p}}\asymp e^{-\ell(y_{n})+\ell(t)-2t-pnt+\psi^{\prime}(y_{n})y_{n}},\quad z\in D(\sigma_{n},\varepsilon|\sigma_{n}|).$
@@
(5.3) $|a_{m,n}|^{p}\asymp\frac{(\psi^{\prime\prime}(y_{n}))^{1/2}}{(\psi^{\prime\prime}(y_{s_{m}}))^{1/2}}e^{\vartheta(m,n)}\frac{|\gamma_{m}-\sigma_{s_{m}}|^{p}}{|\gamma_{m}-\sigma_{n}|^{p}}$
@@
This completes the proof. ∎
@@
(5.4) $\lesssim|z-w|\sum_{n}|a_{n}|\frac{(\psi^{\prime\prime}(y_{n}))^{\frac{1}{2p}}}{|z-\sigma_{n}||w-\sigma_{n}|}e^{\frac{1}{p}(\psi(y_{n})-\ell(y_{n}))}..$
$\bullet$ The case where $1<p<\infty$. Let $q$ the Hölder conjugate of $p$. Using Hölder’s inequality we obtain
@@
If $\Gamma$ is an interpolating sequence, then for every $k$ there exists $f_{k}\in\mathcal{F}^{p}_{\varphi}$ such that
@@
Hence $d_{\log}(\gamma_{k},\gamma_{j})\gtrsim 1$ and hence $\Gamma$ is logarithmically separated. ∎
@@ 6.1. Sufficient conditions
(6.1) $f_{v}(z):=\sum_{n}v_{n}\|L_{\gamma_{n}}\|\frac{G_{\Gamma}(z)}{G^{\prime}_{\Gamma}(\gamma_{n})(z-\gamma_{n})},\quad z\in\mathbb{C}.$
@@
(6.2) $A_{n,m}:=\frac{\|L_{\gamma_{n}}\|}{\|L_{\sigma_{m}}\|}\frac{G_{\Gamma}(\sigma_{m})}{G^{\prime}_{\Gamma}(\gamma_{n})(\sigma_{m}-\gamma_{n})}$
@@
(6.3) $\left|A_{n,m}\right|\lesssim\exp\left(-\left[\frac{1}{2}-\Delta_{N}-\varepsilon\right]|y_{m+j}-y_{n}|\right).$
@@ 6.2. Necessary conditions
(6.4) $|f_{\gamma}(z)|^{p}\lesssim\|f_{\gamma}\|^{p}_{\varphi,p}\|L_{z}\|^{p}\lesssim\frac{\left(\psi^{\prime\prime}(\log|z|)\right)^{\frac{1}{2}}}{\|L_{\gamma}\|^{p}(1+|z|^{2})}e^{p\varphi(z)},\quad z\in\mathbb{C}.$
@@
(6.5) $\left|G_{\Gamma}(\sigma_{m_{k}})\right|\asymp\frac{|\sigma_{m_{k}}-\gamma_{n_{k}}|}{|\gamma_{n_{k}}|}\prod_{j=0}^{n_{k}-1}\frac{\sigma_{m_{k}}}{|\gamma_{j}|},\qquad\left|G^{\prime}_{\Gamma}(\gamma_{n_{k}})\right|\asymp\frac{1}{|\gamma_{n_{k}}|}\prod_{j=0}^{n_{k}-1}\frac{|\gamma_{n_{k}}|}{|\gamma_{j}|}.$
@@
(6.6) $e^{pn_{k}(y_{m_{k}}-x_{n_{k}})}=\left(\frac{\sigma_{m_{k}}}{|\gamma_{n_{k}}|}\right)^{pn_{k}}\asymp\left|f_{\gamma_{n_{k}}}(\sigma_{m_{k}})\right|^{p}\lesssim\frac{\left(\psi^{\prime\prime}(y_{m_{k}})\right)^{\frac{1}{2}}}{\|L_{\gamma_{n_{k}}}\|^{p}}e^{\psi(y_{m_{k}})-2y_{m_{k}}}.$
@@
(6.7) $A_{n,m}:=\frac{\|L_{\gamma_{n}}\|}{\|L_{\sigma_{m}}\|}\frac{G_{\Gamma}(\sigma_{m})}{G^{\prime}_{\Gamma}(\gamma_{n})(\sigma_{m}-\gamma_{n})}$
@@
(6.8) $p\sum_{k=1}^{m}y_{k}=-\psi^{\prime}(y_{m})y_{m}+\psi(y_{m})-\frac{p}{2}y_{m}-\left(\frac{p^{2}}{12}+o(1)\right)\frac{1}{\psi^{\prime\prime}(y_{m})}+O(1).$
@@
(6.9) $|A_{n,m}|^{p}\gtrsim\exp\left[-\frac{p}{2}|y_{n}-y_{m+j}|+C_{\psi}\left(\frac{1}{\psi^{\prime\prime}(y_{n})}-\frac{1}{\psi^{\prime\prime}(y_{m+j})}\right)\pm p\sum_{k=m+j+1}^{n}\delta_{k}\right].$
@@
(7.1) $F_{v}(z)=\sum_{n\geq-1}v_{n}e^{\varphi(\gamma_{n})}\frac{G_{\Gamma^{*}}(z)}{G_{\Gamma^{*}}^{\prime}(\gamma_{n})(z-\gamma_{n})}$
@@
To prove Theorem 1.3, we adopt an approach inspired by [4, 24]. To establish the necessary conditions for a sequence to be sampling or interpolating, we rely on the following lemma. The proof in our setting is identical to that of [4, Lemma 5.1] (see also [21, Lemma 40]), using the point evaluation norm estimates and the canonical product bounds derived above.
@@
(8.1) $\frac{1}{y_{(kN+2)M}-y_{(kN+1)M}}\left|\sum_{n=(kN+1)M+1}^{(kN+N+1)M}\ \delta_{n}\right|\ \leq\ C.$
@@
(8.2) $|\delta_{n}|<\eta(pM(m+1)+2-\frac{p}{2})-\eta(pmM+2-\frac{p}{2})$
@@
(8.5) $|\delta_{n}|<\eta(pM(j+1)+2-\frac{p}{2})-\eta(pjM+2-\frac{p}{2})$
@@
Acknowledgements
The authors would like to thank Sasha Borichev for his helpful remarks and suggestions. The first author was partially supported by ANR Project ANR-24-CE40-5470.